find the slope of (0,1)and(3,0)
Found 2 solutions by josmiceli, lenny460:
Answer by josmiceli(19441)

Slope = ( change in y ) / ( change in x )
slope = ( 0 - 1 ) / ( 3 - 0 )
slope = -1 / 3

Answer by lenny460(1073)

find the slope of (0,1)and(3,0)

The formula for finding the slope of a straight line:
(y2 - y1) / (x2 - x1)
(0 - 1) / (3 - 0)
0 -1 = -1
3 - 0 = 3
The Slope (m) = -1/3
The other way:
(1 - 0) / (0 - 3)
1 - 0 = 1
0 - 3 = -3
The Slope (m) = 1/-3 = -1/3

The Answer:
The Slope (m) = -1/3




Lennox Obuong
Algebra Student
Email: obuong3@aol.com